TiO 2 is made to react with thw stream of hydrogen to produce new oxide Ti x O Y and water. if 1.6 gm of titanimium oxide produces 1.44gm of Ti X O Y. find x and y.

TiOis made to react with thw stream of hydrogen to produce new oxide TixO and water. if 1.6 gm of titanimium oxide produces 1.44gm of TiXOY. find x and y.

Grade:11

1 Answers

Vikas TU
7 years ago
Apply POAC for Ti atom,
1 * 1.6/80 = x * 1.44/(48x + 16y)
1/50 = 1.44x/((48x + 16y)......................................(1)
 
Apply POAC for same O atom and solve the eqns for x and y.